# Understanding LLM Embeddings for Regression

### v2

@@ 2.1 Modeling
To demonstrate consistent results over different families of language models, we benchmark over both the T5 (Raffel et al., 2020) and Gemini 1.0 (Google, 2024) families, which use different architectures (encoder-decoder and decoder-only), different vocabulary sizes (32K and 256K), and embedding dimensions (see Appendix B.2) respectively. However, to remain consistent with the definition of embedding, we follow previous literature (Li et al., 2020; Reimers and Gurevych, 2019) and use average-pooling as the canonical method of aggregating Transformer outputs, and thus the embedding dimension $d_{\text{llm}}$ is equivalent to the the output feature dimension $f$ following a forward pass.
Similar to previous work (Song et al., 2024; Nguyen et al., 2024), for string representations of $x$ from any regression task, by default we use a key-value JSON format with consistent ordering of keys, i.e. {param1:value1,param2:value2,…}, with specific examples shown in Appendix C.
@@ 2.2 Regression Tasks
For regression tasks, we first use synthetic, closed-form objective functions in order to produce controlled studies in which we may query any $x$ from the input space. Our synthetic functions are defined from the standard Black-Box Optimization Benchmarking (BBOB) suite (Elhara et al., 2019). To avoid confounding terminology between embedding "dimension" $d$ and the intrinsic "dimension" of an objective $f$, we denote the latter as "degree-of-freedom" (DOF), and thus $f(\cdot)$ is dependent on input coordinates $x^{(1)},\ldots,x^{(\textrm{DOF})}$, each of which is between $[-5,5]$. This provides a comprehensive variety of both convex and non-convex objective landscapes to regress upon.
@@ 3.1 High Dimensional Regression
We begin by demonstrating cases in which LLM embeddings better represent inputs over high DOF spaces than traditional representations. In Figure 2, we show that for a subset of functions, LLM embeddings possess surprising robustness, retaining the same performance for varying DOFs whereas traditional baselines such as XGBoost and MLPs significantly falter over higher DOFs.
@@ 3.3 Model Effects
In this subsection, we comprehensively investigate the impact of many common LLM factors on regression performance.
Are Larger Models Always Better? Within the research community, the prevailing assumption is that there exists a direct correlation between language model size and performance improvement. However, with the rise of leaderboards such as LMSYS (LMS, 2023), smaller models have been shown to outperform larger competitors, due to differences in their "recipe", such as training data quality, pre-training and post-training techniques, and architecture.
@@
Figure 7: Kendall-Tau regression comparisons when comparing to random initialization (left) and vocabulary embeddings (right). Each bar is averaged across 75 tasks per family.
In Figure 7, we see that the default mode of applying a forward pass of a pre-trained model performs the best, as expected. However, it is worth noting that in some tasks such as AutoML and L2DA, the improvement is surprisingly quite minimal, suggesting that applying forward passes by pretrained models does not always help for regression.
@@
The results of Figures 7 and 8 combined lead to additionally surprising conclusions, such as language-to-numeric transfer. For instance, inputs $x$ from Init2Winit tasks only possess numeric values, and as expected, removing feature names does not significantly change regression results. Yet applying forward passes by pre-trained T5 models still benefits regression, despite the fact that T5’s pre-training data contains mostly web-corpus data which is unlikely to contain significant amounts of scientific or numeric information (Dodge et al., 2021).
@@ B.1 Hyperparameters Used
B.2 Embedding Sizes
@@
Appendix C Example String Representations
@@
BBOB x0:0.32, x1:-4.21, x2:3.12, x3:1.56
AutoML batch_size:128, ml_feature_selection_threshold:0.05, model_type:’DNN_ESTIMATOR’, activation_fn:’selu’, batch_norm:’False’,
bucketization_strategy:’mdl’,dropout:0.071, hidden_units:359

### v3

@@ 2.1 Modeling Specifics
To demonstrate consistent results over different families of language models, we benchmark over both the T5 (Raffel et al., 2020) and Gemini 1.0 (Google, 2024) families, which use different architectures (encoder-decoder and decoder-only), different vocabulary sizes (32K and 256K), and embedding dimensions, respectively. Following (Li et al., 2020; Reimers and Gurevych, 2019), we use average-pooling as the canonical method of aggregating Transformer outputs, as empirical comparisons in Appendix A.4 found this was optimal. Thus the embedding dimension $d_{\text{llm}}$ is equivalent to the the output feature dimension $f$ following a forward pass, and specific sizes of $d_{\text{llm}}$ can be found in Appendix B.3.
Similar to previous work (Song et al., 2024; Nguyen et al., 2024), for string representations of $x$ from any regression task, by default we use a key-value JSON format with consistent ordering of keys, i.e. {param1:value1,param2:value2,…}, with specific examples shown in Appendix C.2.
@@ 2.2 Regression Tasks
Appendix C.1 contains full details on our regression tasks. We first use synthetic, closed-form functions in order to produce controlled studies in which we may query any $x$ from the input space. We use 23 functions defined from the standard Black-Box Optimization Benchmarking (BBOB) suite (Elhara et al., 2019), supporting continuous inputs of any dimension. To avoid confounding terminology between embedding "dimension" $d$ and the intrinsic "dimension" of an objective $f$, we denote the latter as "degree-of-freedom" (DOF), and thus $f(\cdot)$ is dependent on input coordinates $x^{(1)},\ldots,x^{(\textrm{DOF})}$, each of which is between $[-5,5]$. This provides a comprehensive variety of both convex and non-convex objective landscapes to regress upon.
@@ 3.1 High Dimensional Regression
We begin by demonstrating cases in which LLM embeddings better represent inputs over high degree-of-freedom spaces than traditional representations. In Figure 2, we show that for a subset of functions, LLM embeddings possess surprising robustness, retaining the same performance for varying DOFs whereas traditional baselines such as XGBoost and MLPs significantly falter over higher DOFs. Appendix A.5 demonstrates this is inherent to the LLM embeddings, regardless of the regression head used (e.g., XGBoost).
@@ 3.3 Model Effects
In this subsection, we comprehensively investigate the impact of many common LLM factors such as model size and pretraining on regression performance. Additionally, we also ablate the effects of different string representations and LLM subcomponents such as vocabulary embeddings.
Are Larger Models Always Better? Within the research community, the prevailing assumption is that there exists a direct correlation between language model size and performance improvement.
@@
However, with the rise of leaderboards such as LMSYS (LMS, 2023), smaller models have been shown to outperform larger competitors, due to differences in their "recipe", such as training data quality, pre-training and post-training techniques, and architecture.
@@
Figure 7: Kendall-Tau regression comparisons when comparing to random initialization (left) and vocabulary embeddings (right) on BBOB regression tasks. Each bar is averaged across all BBOB functions for a given model size and DOF.
Figure 8: Same as Figure 7, but applied for real world regression tasks. Each bar is averaged across 75 tasks per family.
We first perform our study over BBOB tasks which use purely numeric tokens, where one may question the benefit of language pretraining over English corpus data. Surprisingly, Figure 7 shows that applying forward passes of pretrained models still helps, especially with larger sizes. However, it is worth mentioning that even features from randomly initialized models and vocabulary embeddings alone are also dimensionally robust, as they do not suffer performance drops with higher DOFs. For real-world tasks in Figure 8, it is more expected that applying forward passes by larger pretrained models helps. However, there is more variance in the improvement gaps, as the improvement is surprisingly quite minimal for some tasks such as AutoML and L2DA, while significant for Init2Winit and XLA.
@@
The results of Figures 8 and 9 combined lead to additionally surprising conclusions, such as language-to-numeric transfer. For instance, inputs $x$ from Init2Winit tasks only possess numeric values, and as expected, removing feature names does not significantly change regression results. Yet applying forward passes by pre-trained T5 models still benefits regression, despite the fact that T5’s pre-training data contains mostly web-corpus data which is unlikely to contain significant amounts of scientific or numeric information (Dodge et al., 2021).
@@
4 Discussion: Limitations and Extensions
In this work, our emphasis was to provide more in-depth understanding of LLM embeddings with respect to regression. While we found many different cases in which they outperform traditional representations, we cannot broadly claim that LLM embeddings should always be used in serious applications of regression. Below, we list some limitations of our work and more potential areas of exploration.
Different Modalities and Inputs: Further investigation is needed to understand how LLM embeddings perform with non-tabular data, including combinatorial objects such as graphs, trees, and other complex structures, but also diverse modalities such as images, videos, and audio data.
Prompt Formatting: While we investigated the effects of parameter names for tabular string representations, we did not investigate the effects of different numeric representations. Since LLMs are predominantly pre-trained over human-written text, our $x$ formats also follows, e.g. $1234.5$ is serialized directly into 1234.5. However, these numbers may also be represented using scientific notation (e.g. 1.23e3) or even customized variants, e.g. [1 10e2 2 10e1 3 10e0 4 10e-1 ] as in (Nogueira et al., 2021). We suspect that the fundamental conclusions would remain similar, although the specific numeric results may change.
Different LLM Services: Our work focuses on "depth" of understanding rather than "breadth" of results, although we did find many similar conclusions for both the T5 and Gemini model families, which are quite different in architecture, pre-training data, and post-training. While it remains to be empirically verified, we hypothesize similar conclusions may occur with other model families, such as GPT-4 (OpenAI, 2023), Claude (Anthropic, 2024), and LLaMA (Touvron et al., 2023), especially as they share fundamentally similar approaches with Gemini.
Different Embedding Definitions: As mentioned in the main body, using average-pooling for LLM "embeddings" follows previous literature for consistency, is one of the simplest methods to obtain a fixed dimensional feature vector. and is also better than using max-pooling or last-token logits as shown in Appendix A.4. However, other definitions have been proposed, such as collecting intermediate outputs (Chen et al., 2022). Such additional methods are worth studying in the future for understanding their effects on regression.
In-Context Learning (ICL): An alternative prompting method, particularly natural for decoder-based architectures, would be to place all previous evaluations in the context as "shots" and obtain only the logits from a query $x$. However, this can severely limit the amount of training data allowed, as the context window still has a finite maximum length. In this paper, we primarily focused on the zero-shot case where the context window only contains the query, which allows the downstream MLP to train over unbounded amounts of data.
Computational Costs: Compared to traditional regression techniques which can even be run on CPUs, LLM inference almost always requires accelerator usage, making them more expensive if needed in serious regression tasks. Remote procedure calls to service-based LLMs such as Gemini also adds an additional layer of latency. However, compute costs for inference are orders of magnitude cheaper than for training, and typically only require a few GPUs or TPUs, making embedding-based regression still very feasible for most academic labs or industries.
Smoothness Computations: One limitation is that our smoothness analysis is only feasible when one has online access to $f(\cdot)$ as in the case of BBOB functions but not offline real world data, since one needs to obtain arbitrarily close $(x,x^{\prime})$ pairs to understand local behaviors within the feature space. Our analysis however, may be extendable to any space $\mathcal{X}$ (e.g. combinatorial) which admits a distance metric.
@@ A.3 Performance Correlations
A.4 Pooling Ablations
We ablate the effects of using different pooling mechanisms on T5-XXL logits, and see that average-pooling is consistently the best among the three standard pooling methods (details in Appendix B.2). We further see that using last-token logits not only performs the worst in absolute performance, but also is generally not dimensionally robust in comparison to average-pooling or max-pooling.
Figure 13: Higher ($\uparrow$) is better. Follows the same setting as Figure 2, but comparing different embedding pooling methods.
A.5 XGBoost for LLM Embeddings
In Figure 2 of the main body, XGBoost outperforms MLP when using $\phi_{\text{trad}}$ as input features. However, this pattern does not hold when using $\phi_{\text{LLM}}$, as shown in Figure 14. Here, XGBoost consistently underperforms against an MLP as a regression head for T5 embeddings across all BBOB functions, model sizes, and DOFs.
Despite this difference in head performance, the key observation remains: T5 embeddings exhibit dimensionality robustness and improve with model size regardless of whether XGBoost or an MLP is used as the regression head. This finding suggests that the quality of LLM embeddings is a fundamental property, independent of the specific regression technique employed.
Figure 14: Higher ($\uparrow$) is better. Follows the same setting as Figure 2 but comparing XGBoost and MLP over $\phi_{\text{LLM}}$, when varying BBOB functions (horizontal axis) and T5 model sizes (vertical axis).
@@ B.1 Hyperparameters Used
B.2 Pooling Mechanisms
Tokenized strings may be appended with additional tokens to the maximum sequence length $L$. Thus for example, before a forward pass, a token sequence may actually appear as $[t_{1},t_{2},\ldots,t_{k},1,0,0,0,\ldots]$ where $t_{i}>1$ and 00 and $1$ are padding and optional end-of-sequence (EOS) tokens, respectively. Below, we specify how each pooling mechanism performs postprocessing steps over logits.
  • Average-Pooling (default): Zero-out padding token logits, and average across the entire length axis.
  • Max-Pooling: Same as Average-Pooling, but take the maximum across the length-axis for each coordinate.
  • Last-Token: Only obtain the logit of $t_{k}$ or the EOS token, similar to the <CLS> token method but applicable to decoder-only methods as well.
Note that while Average-Pooling will have its coordinates scaled down by the maximum length $L$, input normalization from the regression head will re-scale them back to valid ranges.
B.3 Embedding Sizes
@@
Appendix C Extended Data Details
C.1 Additional Data Details
BBOB: The Black-Box Optimization Benchmarking (BBOB) suite (Elhara et al., 2019) consists of 23 synthetic objectives which supports evaluation over continuous inputs of any DOF. Such functions comprehensively cover different landscape conditions (e.g. separability and optimality). Every function $f(x)$ by default has a global minimum $f(x)=0$ at exactly the zero-vector input. Example of such functions include:
  • Sphere: $f(x)=\sum_{i=1}^{DOF}(x^{(i)})^{2}$
  • Rastrigin: $f(x)=10\left(DOF-\sum_{i=1}^{DOF}\cos(2\pi x^{(i)})\right)+\lVert x\rVert^{2}$
  • BentCigar: $f(x)=(x^{(1)})^{2}+10^{6}\sum_{i=2}^{DOF}(x^{(i)})^{2}$
Real World Data: The Google Vizier database contains an extensive set of hyperparameter and blackbox optimization objectives, which possess varied landscapes suitable for regression benchmarking.
  • AutoML: Objectives collected from the Vertex AI platform for automated ML model selection and training for tabular or text data. For tabular data, Vertex AI searches over a tree of model and optimizer types, their hyperparameters, data transformation, and other components in the ML pipeline. For text, Vertex AI trains an ML model to classify text data, extract information, or understand the sentiment of the authors. For more information, see: https://cloud.google.com/vertex-ai?#train-models-with-minimal-ml-expertise.
  • Init2Winit: Data from running deterministic, scalable, and well-documented deep learning experiments, with a particular emphasis on optimization and tuning experiments (e.g. ResNets on CIFAR10, Transformers on LM1B). Public codebase can be found in https://github.com/google/init2winit.
  • XLA: Tuning LLM latencies over a variety of different settings for the XLA compiler. Such hyperparameters affect sharding, partitioning, memory consumption, and data limits when placing LLM models on accelerators. Such hyperparameters can be found in OpenXLA https://github.com/openxla/xla.
  • L2DA: Hardware-specific hyperparameters such as memory depth and input-output bandwidths, used to optimize computer hardware performance or reduce costs. Many such similar hyperparameters can be found in https://github.com/srivatsankrishnan/oss-arch-gym.
C.2 String Representations
@@
BBOB [0.3221, -4.2113, 3.1212, 1.5621]
AutoML batch_size:128, ml_feature_selection_threshold:0.05, model_type:’DNN_ESTIMATOR’,
activation_fn:’selu’, batch_norm:’False’, bucketization_strategy:’mdl’,
dropout:0.071, hidden_units:359